OpenStudy (anonymous):

I've been working on this set of equations for about a month; interesting problem. It's an optimization problem. First, we have $$\phi (x) = ((x-a)^2 + (y-b)^2 + (z-c)^2)^{1/2}$$

OpenStudy (anonymous):

Now, the second derivative of this is: $$\frac{\partial^2 \phi}{\partial x^2} = \frac{b^2 + c^2 - 2by + y^2 -2cz + z^2}{((a-x)^2 + (b-y)^2 + (c-z)^2)^{3/2}}$$

OpenStudy (anonymous):

Now I have to solve such that: $$\frac{\partial^2 \phi}{\partial x^2} < 0$$

OpenStudy (anonymous):

In the actual problem, we're given \(x, y\) and \(z\), so treat them as real constants.

OpenStudy (anonymous):

Oh yes, sorry. \(\phi\) is actually a function of x, y and z.

OpenStudy (anonymous):

And we want to solve the inequality on a boundary \([\Delta x, \Delta y\, \Delta z)]\)

OpenStudy (primeralph):

So, what now?

OpenStudy (primeralph):

You said y, z are given, right?

OpenStudy (anonymous):

x, y, and z are given. I need to solve for each variable.

OpenStudy (primeralph):

What are the variables?

OpenStudy (anonymous):

a, b and c.

OpenStudy (anonymous):

Whether you consider a, b, c constants or x, y and z constants doesn't matter. Easily interchangeable. It may be more convenient to solve for x, y and z as variables.

OpenStudy (primeralph):

You might want to look at this as a growing (or shrinking) sphere with the center changing where phi is the radius. Or simply take partials in y, z to observe their minima and use that.

OpenStudy (anonymous):

We're observing the maxima.

OpenStudy (anonymous):

We're trying to find the maximal distance between two points on a boundary.

OpenStudy (anonymous):

We might be able to set up a differential equation or something with the growing/shrinking sphere visualization.

OpenStudy (dan815):